# On orthogonal decompositions of hermitian Higgs bundles

### v1

@@ 2. Hermitian Higgs bundles and the second fundamental form
The aim of this section, is to revisit a couple of results on hermitian Higgs bundles previously known in literature [2] and which are extensions of propositions 2 and 3. We include here alternative proofs of these results without using local computations. In this section we also recall relevant information about Higgs bundles and fix the remaining part of the terminology that will be used throughout the article. In order to do this, we begin here reviewing some standard definitions on Higgs bundles (for more details the reader can see [5, 6] or the pioneering work [15]).
An hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle [15] and $(E,h)$ is a holomorphic hermitian vector bundle [12]. In particular, the Higgs field $\Phi\in A^{1,0}({\rm End}E)$ is holomorphic and satisfies $\Phi\wedge\Phi=0$. Using the hermitian metric $h$ one has an hermitian conjugate ${\bar{\Phi}}_{h}\in A^{0,1}({\rm End}E)$ of the Higgs field and we get in the Higgs bundles setting the following connection:
(2.1) ${\cal D}_{h}=D_{h}+\Phi+{\bar{\Phi}}_{h}\,,$
usually known as the Hitchin-Simpson connection of $({\mathfrak{E}},h)$ [2]. A straightforward computation shows that its curvature ${\cal R}_{h}={\cal D}_{h}\wedge{\cal D}_{h}$ is given by
(2.2) ${\cal R}_{h}=R_{h}+D^{\prime}_{h}\phi+d^{\prime\prime}{\bar{\Phi}}_{h}+[\Phi,{\bar{\Phi}}_{h}]\,.$
@@
where $I$ denotes here the identity morphism of $\Omega^{1,0}$, i.e., it is a morphism of the corresponding holomorphic vector bundles in which the obvious square diagram commutes. If ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle, a Higgs subbundle ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a $\Phi$-invariant holomorphic subbundle $S\subset E$. Hence, if $\Phi_{S}$ denotes the restriction of the Higgs field $\Phi$ to $S$, the pair $(S,\Phi_{S})$ is itself a Higgs bundle and the inclusion induces a morphism of Higgs bundles ${\mathfrak{S}}\longrightarrow{\mathfrak{E}}\,$.
The following is a well known result on Higgs bundles, however it plays a fundamental role in this article and hence we include it explicitly.
@@
Let $({\mathfrak{E}},h)$ be an hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ a Higgs subbundle. Since $S$ is $\Phi$-invariant it follows that $S^{\bot}$ is ${\bar{\Phi}}_{h}$-invariant. [^2] Let $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\bot})$. From the definition of hermitian conjugate and the $\Phi$-invariance of $S$ we get $h(\bar{\Phi}_{h}\eta,\xi)=h(\eta,\Phi\xi)=0$ and $\bar{\Phi}_{h}\eta\in A^{1}(S^{\bot})\,$. Moreover, from Section 1 we know that $Q\cong S^{\bot}$ (as $C^{\infty}$ complex vector bundles) and hence using Lemma 1 we get that $S^{\bot}$ is $\Phi$-invariant, which in turn implies that $S$ is $\bar{\Phi}_{h}$-invariant. From the above we immediately get the following result.
Lemma 2.
If ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, then the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant if and only if it is ${\cal D}_{h}$-invariant.
For future reference, we introduce the following terminology. If ${\mathfrak{S}}$ is a Higgs subbundle of ${\mathfrak{E}}$ and the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant, we say that the Higgs subbundle $\mathfrak{S}$ is $D_{h}$-invariant. A similar definition applies to the Hitchin-Simpson connection ${\cal D}_{h}$.
If now $\xi\in A^{0}(S)$ and we use (2.1) and (1.3) we get a decomposition
${\cal D}_{h}\xi=(D_{h,S}+\Phi+\bar{\Phi}_{h})\xi+A_{h}\xi\,,$
with $(D_{h,S}+\Phi+\bar{\Phi}_{h})\xi\in A^{1}(S)$ and $A_{h}\xi\in A^{1}(S^{\bot})$. Notice that here we are using not only that $S$ is $\Phi$-invariant but also $\bar{\Phi}_{h}$-invariant. By defining
(2.4) ${\cal D}_{h,S}=D_{h,S}+(\Phi+\bar{\Phi}_{h})\lvert_{S}\,$
@@
Let $({\mathfrak{E}},h)$ be an hermitian Higgs bundle, where ${\cal D}_{h}$ is the Hitchin-Simpson connection of it and let ${\mathfrak{S}}\subset{\mathfrak{E}}$ be a Higgs subbundle. Then the operator ${\cal D}_{h,S}$ defined in (2.4) is the Hitchin-Simpson connection of $({\mathfrak{S}},h_{S})$ and $A_{h}$ is the second fundamental form of Proposition 2, i.e., ${\cal D}_{h,S}={\cal D}_{h_{S}}$ and $A_{h}\in A^{1,0}({\rm Hom}(S,S^{\bot}))$.
This result can be seen as an extension to Higgs bundles of the classical Proposition 2. We would like to emphasize that, strictly speaking, Proposition 5 is already known in literature and it appears in the form of Proposition 2.3 in [2] (parts (i) and (ii)). However, in that reference the result is found and proved in a different way, in particular in that reference the authors introduce the main objects using local unitary frame fields. Here we only make explicit that Proposition 5 is closely related to a well known result in holomorphic hermitian vector bundles and that it can be also obtained without using a local computation.
@@
${\cal D}_{h}\eta=B_{h}\eta+(D_{h,S^{\bot}}+\Phi+\bar{\Phi}_{h})\eta\,,$
with $(D_{h,S}+\Phi+\bar{\Phi}_{h})\eta\in A^{1}(S^{\bot})$ and $B_{h}\eta\in A^{1}(S)$. Notice that this time we are using that $S^{\bot}$ is $\Phi$-invariant and also $\bar{\Phi}_{h}$-invariant. By defining
(2.6) ${\cal D}_{h,S^{\bot}}=D_{h,S^{\bot}}+(\Phi+\bar{\Phi}_{h})\lvert_{S^{\bot}}\,$
@@
(2.8) ${\cal D}_{h,S^{\bot}}\equiv{\cal D}_{h,Q}:A^{0}(Q)\longrightarrow A^{1}(Q)$
@@
Let $({\mathfrak{E}},h)$, ${\cal D}_{h}$ and ${\mathfrak{S}}$ as in Proposition 5 and ${\mathfrak{Q}}$ the quotient Higgs bundle given by Lemma 1. Then the operator ${\cal D}_{h,Q}$ defined in (2.8) is the Hitchin-Simpson connection of $({\mathfrak{Q}},h_{Q})$ and $B_{h}$ is the same form of Proposition 3, i.e., ${\cal D}_{h,Q}={\cal D}_{h_{Q}}$ and $B_{h}\in A^{0,1}({\rm Hom}(S^{\bot},S))$.
The above proposition can be seen as a Higgs extension of the classical Proposition 3. Notice that the form $B_{h}$ is mentioned as part (iii) of Proposition 2.3 in [2]. The remaining content of this result is already implicit in the local computations in that reference. Again, here we only make explicit that Proposition 6 is closely related to a well known result in holomorphic hermitian vector bundles and that it can be obtained without using a local computation.
We would like to emphasize that the forms $A_{h}$ and $B_{h}$ of propositions 5 and 6 are the same as the classical forms for holomorphic subbundles of hermitian holomorphic bundles, a fact also mentioned in [2]. Therefore, if $({\mathfrak{E}},h)$ is an hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, the second fundamental form of ${\mathfrak{S}}$ in $({\mathfrak{E}},h)$ is just the second fundamental form of the holomorphic subbundle $S$ in $(E,h)$. Now, as in the case of holomorphic hermitian vector bundles, (2.5) and (2.7) can be extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. Therefore, for any $\xi\in A^{0}(E)$ we get
${\cal R}_{h}\xi=({\cal R}_{h,S}+B_{h}\wedge A_{h}+{\cal D}_{h}A_{h})\xi\,,$
@@
${\cal R}_{h}\eta=({\cal R}_{h,S^{\bot}}+A_{h}\wedge B_{h}+{\cal D}_{h}B_{h})\eta\,.$
As it is usual, the above expressions can be used to rewrite ${\cal R}_{h}$ in a matrix form as:
(2.9) $\mathcal{R}_{h}=\begin{pmatrix}\mathcal{R}_{h,S}+B_{h}\wedge A_{h}&\mathcal{D}_{h}B_{h}\\ \mathcal{D}_{h}A_{h}&\mathcal{R}_{h,S^{\bot}}+A_{h}\wedge B_{h}\end{pmatrix},$
which is the Higgs extension of the Chern curvature matrix formula (1.7). The expression (2.9) can be considered as the Gauss-Codazzi type equations for hermitian Higgs bundles (cf. [2], p. 7).

### v2

@@ 3. Hermitian Higgs bundles and the second fundamental form
The aim of this section, is to revisit a couple of results on hermitian Higgs bundles previously known in literature [2]. Here, we show that such results admit generalizations that can be viewed as natural extensions of propositions 2 and 3. We include here alternative proofs of these results without using local computations. In this section we also recall relevant information about Higgs bundles and fix the remaining part of the terminology that will be used throughout the article. In order to do this, we begin here reviewing some standard definitions on Higgs bundles (for more details the reader can see [6, 7] or the pioneering work [17]).
A hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle [17] and $(E,h)$ is a holomorphic hermitian vector bundle [13]. In particular, the Higgs field $\Phi\in A^{1,0}({\rm End}E)$ is holomorphic and satisfies $\Phi\wedge\Phi=0$. Using the hermitian metric $h$ one has a hermitian conjugate ${\bar{\Phi}}_{h}\in A^{0,1}({\rm End}E)$ of the Higgs field and we get in the Higgs bundles setting the following connection:
(3.1) ${\mathcal{D}}_{h}=D_{h}+\Phi+{\bar{\Phi}}_{h}\,$,
usually known as the Hitchin-Simpson connection of $({\mathfrak{E}},h)$ [2]. A straightforward computation shows that its curvature ${\mathcal{R}}_{h}={\mathcal{D}}_{h}\wedge{\mathcal{D}}_{h}$ is given by
(3.2) ${\mathcal{R}}_{h}=R_{h}+D^{\prime}_{h}\Phi+d^{\prime\prime}{\bar{\Phi}}_{h}+[\Phi,{\bar{\Phi}}_{h}]\,.$
@@
where $I$ denotes here the identity morphism of $\Omega^{1,0}$, i.e., it is a morphism of the corresponding holomorphic vector bundles in which the obvious square diagram commutes. If ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle, a Higgs subbundle ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a $\Phi$-invariant holomorphic subbundle $S\subset E$. Hence, if $\Phi\lvert_{S}$ is the restriction of the Higgs field $\Phi$ to $S$, the pair $(S,\Phi\lvert_{S})$ is itself a Higgs bundle and the inclusion induces a morphism of Higgs bundles ${\mathfrak{S}}\longrightarrow{\mathfrak{E}}\,$. Whenever a metric $h_{S}$ is defined, we simply write ${\bar{\Phi}}_{h_{S}}$ to denote the hermitian conjugate of $\Phi\lvert_{S}$ with respect to $h_{S}$. If $\mathfrak{S}$ is a Higgs subbundle and $P_{S}$ and $P_{S^{\perp}}$ are the standard projections onto $S$ and $S^{\perp}$, we denote by $P_{S}\Phi$ and $P_{S^{\perp}}\Phi$ the compositions of $P_{S}$ and $P_{S^{\perp}}$ with $\Phi$. A similar definition applies to ${\bar{\Phi}}_{h}$.
The following is a well-known result on Higgs bundles, however it plays a fundamental role in this article and hence we include it explicitly.
@@
Let $({\mathfrak{E}},h)$ be a hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ a Higgs subbundle. Since $S$ is $\Phi$-invariant it follows that $S^{\bot}$ is ${\bar{\Phi}}_{h}$-invariant. In fact, if $\eta\in A^{0}(S^{\bot})$ and $\xi\in A^{0}(S)$, then we get $h(\bar{\Phi}_{h}\eta,\xi)=h(\eta,\Phi\xi)=0$ and $\bar{\Phi}_{h}\eta\in A^{1}(S^{\bot})\,$. Moreover, in general $S^{\perp}$ is not $\Phi$-invariant [^5] Let $M$ be an $1$-dimensional compact complex manifold of genus $1$, let $\mathcal{O}_{M}$ be its structure sheaf and choose a nowhere-vanishing $1$-form $\varphi$. Set $E=\mathcal{O}_{M}\oplus\mathcal{O}_{M}$ and define $\Phi\in A^{1,0}(\mathrm{End}E)$ by $\Phi=\begin{pmatrix}0&\varphi\\ 0&0\end{pmatrix}$. The pair $\mathfrak{E}=(E,\Phi)$ is a rank $2$ Higgs bundle over $M$. Let $h$ be the product hermitian metric on $E$ and consider a holomorphic orthogonal frame $\{s_{1},s_{2}\}$, so the subbundles $S=\langle s_{1}\rangle$ and $S^{\perp}=\langle s_{2}\rangle$ give an orthogonal decomposition $E=S\oplus S^{\perp}$ with $A_{h}=0$ and such that $\Phi s_{1}=0$ and $\Phi s_{2}=\varphi s_{1}$. Thus $S$ is $\Phi$-invariant but $S^{\perp}$ is not. Consequently, from Proposition 4 the decomposition $E=S\oplus S^{\perp}$ is also holomorphic and the pair $\mathfrak{S}=(S,\Phi\lvert_{S})$ is a Higgs subbundle (see [10] for more details on this type of constructions). and, as we will see in Section 4, this fact plays a crucial role in the decomposition of a Higgs bundle in terms of Higgs subbundles. If $S^{\perp}$ is not $\Phi$-invariant, a minor modification of the previous arguments implies that also $S$ is not ${\bar{\Phi}}_{h}$-invariant. [^6] In fact, if $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\perp})$ it follows that $0\neq h(\xi,\Phi\eta)=h({\bar{\Phi}}_{h}\xi,\eta)$ and hence, in general, ${\bar{\Phi}}_{h}\xi$ has a component in $A^{1}(S^{\perp})$. Consequently, given any section $\xi\in A^{0}(S)$ we have
(3.4) ${\bar{\Phi}}_{h}\xi=P_{S}{\bar{\Phi}}_{h}\xi+P_{S^{\perp}}{\bar{\Phi}}_{h}\xi.$
For future reference, we introduce the following terminology. If $\mathfrak{S}$} is a Higgs subbundle of $\mathfrak{E}$} and the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant, we say that the Higgs subbundle $\mathfrak{S}$ is $D_{h}$-invariant. A similar definition applies to the Hitchin-Simpson connection ${\mathcal{D}}_{h}$.
If now $\xi\in A^{0}(S)$ and we use (2.6), (3.1) and (3.4) we get a decomposition
${\mathcal{D}}_{h}\xi=(D_{h,S}+\Phi+P_{S}{\bar{\Phi}}_{h})\xi+(A_{h}+P_{S^{\perp}}{\bar{\Phi}}_{h})\xi\,$,
with $(D_{h,S}+\Phi+P_{S}{\bar{\Phi}}_{h})\xi\in A^{1}(S)$ and $(A_{h}+P_{S^{\perp}}{\bar{\Phi}}_{h})\xi\in A^{1}(S^{\bot})$. One observes that the $\mathrm{End}S$-valued form $P_{S}{\bar{\Phi}}_{h}\lvert_{S}:S\longrightarrow S\otimes\Omega^{0,1}$ is just the hermitian conjugate of $\Phi\lvert_{S}$ with respect to $h_{S}$. [^7] If $\xi,\eta\in A^{0}(S)$ then $h_{S}({\bar{\Phi}}_{h_{S}}\xi,\eta)=h_{S}(\xi,\Phi\eta)=h(\xi,\Phi\eta)=h({\bar{\Phi}}_{h}\xi,\eta)=h(P_{S}{\bar{\Phi}}_{h}\xi,\eta)=h_{S}(P_{S}{\bar{\Phi}}_{h}\xi,\eta),$ where in the penultimate equality we have used (3.4). Therefore ${\bar{\Phi}}_{h_{S}}=P_{S}{\bar{\Phi}}_{h}\lvert_{S}$. By defining
(3.5) ${\mathcal{D}}_{h,S}=D_{h,S}+\Phi\lvert_{S}+P_{S}{\bar{\Phi}}_{h}\lvert_{S}\,$
and
(3.6) ${\mathcal{A}}_{h}=A_{h}+P_{S^{\perp}}{\bar{\Phi}}_{h}\lvert_{S}$
@@
Let $({\mathfrak{E}},h)$ be a hermitian Higgs bundle and ${\mathcal{D}}_{h}$ its Hitchin-Simpson connection. Let ${\mathfrak{S}}\subset{\mathfrak{E}}$ be a Higgs subbundle. Then the operator ${\mathcal{D}}_{h,S}$ defined in (3.5) is the Hitchin-Simpson connection of $({\mathfrak{S}},h_{S})$ and ${\mathcal{A}}_{h}$ is a $1$-form with values in $\mathrm{Hom}(S,S^{\perp})$, i.e., ${\mathcal{D}}_{h,S}={\mathcal{D}}_{h_{S}}$ and ${\mathcal{A}}_{h}\in A^{1}({\rm Hom}(S,S^{\bot}))$.
This result can be seen as an extension to Higgs bundles of Proposition 2. We would like to emphasize that, strictly speaking, Proposition 5 is also a generalization of Proposition 2.3 (parts (i) and (ii)) in [2]. Moreover, their result is established and proved using a different approach. In particular, in that work the authors introduce the main objects using local unitary frame fields. Here we make explicit that Proposition 5 is closely related to a well-known result in holomorphic hermitian vector bundles and that it can be obtained without using local computations. Since in general $S^{\perp}$ is not $\Phi$-invariant, given any section $\eta\in A^{0}(S^{\perp})$ we have
(3.8) $\Phi\eta=P_{S^{\perp}}\Phi\eta+P_{S}\Phi\eta.$
@@
${\mathcal{D}}_{h}\eta=(B_{h}+P_{S}\Phi)\eta+(D_{h,S^{\bot}}+P_{S^{\perp}}\Phi+\bar{\Phi}_{h})\eta\,$,
with $(D_{h,S^{\perp}}+P_{S^{\perp}}\Phi+\bar{\Phi}_{h})\eta\in A^{1}(S^{\bot})$ and $(B_{h}+P_{S}\Phi)\eta\in A^{1}(S)$. Notice that this time we are using that $S^{\bot}$ is $\bar{\Phi}_{h}$-invariant and the $\mathrm{End}S^{\perp}$-valued form ${\bar{\Phi}}_{h}\lvert_{S^{\perp}}:S^{\perp}\longrightarrow S^{\perp}\otimes\Omega^{0,1}$ is the hermitian conjugate of $P_{S^{\perp}}\Phi\lvert_{S^{\perp}}:S^{\perp}\longrightarrow S^{\perp}\otimes\Omega^{1,0}$ with respect to $h_{S^{\perp}}$. [^8] In fact, if $\xi,\eta\in A^{0}(S^{\perp})$ we obtain $h_{S^{\perp}}((\overline{P_{S^{\perp}}\Phi})_{h_{S^{\perp}}}\xi,\eta)=h_{S^{\perp}}(\xi,P_{S^{\perp}}\Phi\eta)=h(\xi,P_{S^{\perp}}\Phi\eta)=h(\xi,\Phi\eta)=h({\bar{\Phi}}_{h}\xi,\eta)=h_{S^{\perp}}({\bar{\Phi}}_{h}\xi,\eta),$ where the third equality follows from (3.8). Consequently $(\overline{P_{S^{\perp}}\Phi})_{h_{S^{\perp}}}={\bar{\Phi}}_{h}\lvert_{S^{\perp}}$. By defining
(3.9) ${\mathcal{D}}_{h,S^{\bot}}=D_{h,S^{\bot}}+P_{S^{\perp}}\Phi\lvert_{S^{\perp}}+\bar{\Phi}_{h}\lvert_{S^{\perp}}\,$
and
(3.10) ${\mathcal{B}}_{h}=B_{h}+P_{S}\Phi\lvert_{S^{\perp}},$
@@
(3.12) ${\mathcal{D}}_{h,S^{\bot}}\equiv{\mathcal{D}}_{h,Q}:A^{0}(Q)\longrightarrow A^{1}(Q)$
@@
Let $({\mathfrak{E}},h)$, ${\mathcal{D}}_{h}$ and $\mathfrak{S}$} as in Proposition 5 and $\mathfrak{Q}$} the quotient Higgs bundle given by Lemma 1. Then the operator ${\mathcal{D}}_{h,Q}$ defined in (3.12) is the Hitchin-Simpson connection of $({\mathfrak{Q}},h_{Q})$ and ${\mathcal{B}}_{h}$ is a $1$-form with values in ${\rm Hom}(S^{\bot},S)$, i.e., ${\mathcal{D}}_{h,Q}={\mathcal{D}}_{h_{Q}}$ and ${\mathcal{B}}_{h}\in A^{1}({\rm Hom}(S^{\bot},S))$.
The above proposition is a Higgs extension of the classical Proposition 3. Again, here we make explicit that Proposition 6 is closely related to a well-known result in holomorphic hermitian vector bundles and that it can be obtained without using local computations. In general, the forms $\mathcal{A}_{h}$ and $\mathcal{B}_{h}$ in propositions 5 and 6 are not equal to the classical forms $A_{h}$ and $B_{h}$ of the corresponding holomorphic objects. In fact, the operators $P_{S^{\perp}}{\bar{\Phi}}_{h}\lvert_{S}$ and $P_{S}\Phi\lvert_{S^{\perp}}$ in (3.6) and (3.10) represent the obstructions to these equalities. To be more precise, if $({\mathfrak{E}},h)$ is a hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, the forms ${\mathcal{A}}_{h}$ and ${\mathcal{B}}_{h}$ of $\mathfrak{S}$} in $({\mathfrak{E}},h)$ are not equal to the classical forms $A_{h}$ and $B_{h}$ of $S$ in $(E,h)$ unless additional conditions are imposed. Now, as in the case of holomorphic hermitian vector bundles, (3.7) and (3.11) can be extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. Therefore, for any $\xi\in A^{0}(S)$ we get
${\mathcal{R}}_{h}\xi=({\mathcal{R}}_{h,S}+{\mathcal{B}}_{h}\wedge{\mathcal{A}}_{h}+{\mathcal{D}}_{h}{\mathcal{A}}_{h})\xi\,$,
@@
${\mathcal{R}}_{h}\eta=({\mathcal{R}}_{h,S^{\bot}}+{\mathcal{A}}_{h}\wedge{\mathcal{B}}_{h}+{\mathcal{D}}_{h}{\mathcal{B}}_{h})\eta\,.$
As it is usual, the above expressions can be used to rewrite ${\mathcal{R}}_{h}$ in a matrix form as:
(3.13) $\mathcal{R}_{h}=\begin{pmatrix}\mathcal{R}_{h,S}+{\mathcal{B}}_{h}\wedge{\mathcal{A}}_{h}&\mathcal{D}_{h}{\mathcal{B}}_{h}\\ \mathcal{D}_{h}{\mathcal{A}}_{h}&\mathcal{R}_{h,S^{\bot}}+{\mathcal{A}}_{h}\wedge{\mathcal{B}}_{h}\end{pmatrix}$,
which is the Higgs extension of the Chern curvature matrix formula (2.10). The expression (3.13) can be considered as the Gauss-Codazzi type equations for hermitian Higgs bundles.